OpenStudy (anonymous):

The sum of 3 consecutive integers is 60. What is the value of the third integer? (Be sure to type only the third integer in the answer blank. Do not use any variables or words.)

OpenStudy (anonymous):

60 --- = 20 3 since you're adding 3 consecutive integers, they will be around 20

OpenStudy (anonymous):

consider 3 integers as x,x+1,x+2 ....add them to get x as 19 then 3rd integer is 19+2=21

OpenStudy (anonymous):

lets test 20+21+22 = 63 thats not wat we want lets test 19+20+21 = 60 so Yes, these are the numbers !

OpenStudy (anonymous):

oh -.- i already put in the answer.,.lol

OpenStudy (anonymous):

omg ! did u put 20 or what ? answer is 21 btw :o